OpenStudy (anonymous):

The following function represents the profit P(n), in dollars, that a concert promoter makes by selling tickets for n dollars each: P(n) = -250n2 + 3,250n - 9,000 Part A: What are the zeroes of the above function, and what do they represent? Show your work. (4 points)

OpenStudy (cj49):

P(n)=-250n2+3,250n-9,000 0=-250n2+3,250-9,000 0=-250(n2+13n+36) 0=n2+13n+36 0=n-9 and n-4 n=9 and 4